Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

night-report-2015-03-28

Observer : Heather Campbell, Nic Walton
Support team : Southampton (
Szymon Prajs)
TAT :

NOTE: Data transfer problems due to issues with the connection between the observing and data reduction machines (roll over problem from earlier powercuts). Support staff not able to fix. Hence was not possible to analyse the data during the night.  Thus Atel will need to be done next night for the classification targets below.  not fixed but now have a work around so can access the data.

Start of night: some clouds on the horizon

End of night: 

Seeing: Measured ??? at start of night, ??? for most of night.

Afternoon calibrations:  EFOSC arcs and specflats

22.52    Sunset 

23.05 Skyflats UVi

23.23 Standard LTT3864
image analysis

23.55 OGLE173 1h G11+G16

01.00 PS15br griz+G13

01.45 PS15ql classification  extended exposure time by 5 mins to 1200s (not sure if header will have this in) as looked v faint. Ia ~1week after peak, z = 0.098

02.15 LSQ15kp  griz imaging

02.40  CHANGE to SOFI

image analysis seeing about 0.9 in SOFI

02.50 SN2015F - imaging JHK and the BG+RG

04.00 telluric Hip044

04.10 PS15ae - imaging JHK 

05.30 PSN10544 - imaging JHK 

06.00 LSQkp - imaging JHK

07.30 SNhunt281 BG +RG 

10.00 +telluric ~5mins Hip 071559

10.05 Flux standard for SOFI - LTT7989 10mins





Comments